# Semantic Scene Completion Based 3D Traversability Estimation for Off-Road Terrains

### v1

@@ III-C Architecture of ORD-BKI
The research presented in[59] introduces the S-BKI model, which infers grid attributes for discrete semantic grid maps to achieve continuous and dense semantic mapping. To enable a fair comparison, the S-BKI model is built upon, utilizing LiDAR point clouds as input. Based on the vehicle’s geometric passability, a feasible domain recognition model, ORD-BKI, is developed, which employs Bayesian Kernel Inference for semantic scene completion, as illustrated in Fig. 6.
The image semantic segmentation results are projected onto the point clouds to obtain semantically enriched point clouds. Due to the discrete and sparse nature of LiDAR point clouds, it is essential to model the distribution of grid semantic attributes during voxelization to infer the semantic attributes of neighboring grids[59]. The ORD-BKI model employs Bayesian Kernel Inference for semantic scene completion, approximating a Gaussian process, which reduces the time complexity from cubic to linear complexity[61]. The Nearest Neighbor algorithm is then used to assign semantics, generating dense semantic occupancy results. The model utilizes an anisotropic kernel function for inference and analyzes the model results under historical frame point cloud inputs. The geometric passability analysis method employed by the model is consistent with the discussions presented earlier in the article.
@@ IV-A Quantitative Analysis
Experiments are conducted on the RELLIS-OCC dataset for the tasks of 3D semantic occupancy prediction and 3D traversability cost estimation, with the results compared to those obtained by the algorithms presented in[62, 63, 64, 59]. Due to GPU memory constraints, the Monoscene model is trained on $4\times V100\ GPUs$, while the other models are trained on $4\times 3090\ GPUs$. A consistent batch size of 4 is maintained, and each model is trained for 20 epochs.
Evaluation Metrics: The RELLIS-OCC dataset comprises both semantic labels and traversability cost annotations. The traversability cost estimation task is framed as a semantic occupancy prediction task, encompassing four categories: lethal, medium-cost, low-cost, and free. Consequently, traditional metrics employed in 3D semantic occupancy prediction tasks are utilized for evaluation. For the scene completion (SC) task, the Intersection over Union (IoU) of occupied voxels, disregarding their semantic classes, is used as the evaluation metric. For the semantic scene completion (SSC) task, the IoU of each semantic class and the mean IoU (mIoU) across all semantic classes are adopted as evaluation metrics.
The IoU is defined as:
$\text{IoU}=\frac{\text{TP}}{\text{TP}+\text{FP}+\text{FN}}.$ (21)
The mIoU across $C$ classes is calculated as:
$\text{mIoU}=\frac{1}{C}\sum_{i=1}^{C}\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FP}_{i}+\text{FN}_{i}},$ (22)
where TP, FP, and FN represent the number of true positive, false positive, and false negative predictions, respectively, and $C$ is the number of classes.
IV-A1 3D Semantic Occupancy Prediction
The results of the 3D semantic occupancy prediction task are presented in Table II. To ensure a fair comparison, all methods are evaluated using consistent perception ranges and voxel sizes. Fig. 4 illustrates the distribution statistics for voxel labels within the dataset. The limited availability of ground truth training data for categories such as person and object poses significant challenges for many models in accurately recognizing these classes. Additionally, the RELLIS-3D dataset, which serves as the foundation for RELLIS-OCC dataset, was collected in a campus environment with very limited overlap between sequences. This, combined with the rarity of hard-surface scenes in the validation set, leads to generally lower recognition metrics for the hard-surface category.
TABLE II: Monocular Semantic Scene Completion Results on the RELLIS-OCC Dataset.
@@
Grass Tree Hard-surface Object Bush Water Person Mud Rubble mIoU IoU
Monoscene[62] 27.99 11.12 0.000 1.04 18.11 3.28 0.001 3.33 16.50 8.14 33.15
SSCNet[63] 22.74 33.92 0.000 6.59 17.22 13.71 0.000 0.000 26.41 12.06 33.99
SSCNet-full[63] 30.19 33.49 0.000 4.22 19.29 19.30 0.001 0.001 29.5 13.61 38.64
LMSCNet[64] 16.34 33.12 0.000 0.000 19.33 0.000 0.000 0.000 14.31 8.31 31.48
LMSCNet-SS[64] 8.94 36.13 0.000 0.000 18.74 0.000 0.000 0.000 0.000 6.38 33.22
ORDformer(Ours) 25.47 37.09 0.008 2.05 23.65 3.81 0.006 10.34 2.59 10.50 50.20
Despite these challenges, ORDformer demonstrates superior performance across a wide range of categories. Notably, it achieves state-of-the-art results in identifying tree and bush, with IoU scores of 37.09 and 23.65, respectively. Furthermore, while many models struggle with challenging classes such as mud, ORDformer achieves an IoU of 10.34, significantly higher than competing approaches. Most impressively, ORDformer outperforms other models by over 20% in the scene completion (SC) IoU metric, achieving a score of 50.20 compared to the second-best result of 38.64 from SSCNet-full. These results underscore the effectiveness of ORDformer in integrating multimodal data and performing robust occupancy predictions in complex environments.
IV-A2 3D Traversability Cost Estimation
The results of the 3D Traversability Cost Estimation task are presented in Table III. Across all methods, those based on semantic scene completion outperform the ORD-BKI method, which relies on Bayesian Kernel Inference. This underscores the effectiveness of semantic scene completion approaches in enhancing off-road traversable area recognition. Notably, methods using only LiDAR point clouds as input, such as LMSCNet and SSCNet, exhibit inferior performance in the traversability cost estimation task compared to models that incorporate camera input. This suggests that image input is indispensable for accurate traversability cost estimation.
TABLE III: Monocular Traversability Estimate Results on the RELLIS-OCC Dataset.(M-cost: Medium-cost, L-Cost: Low-cost)
Method SSC SSC SC
Lethal M-cost L-cost Free mIoU IoU
@@
ORDformer(Ours) 19.98 13.89 23.50 5.41 15.70 49.78
Our investigation reveals that while methods employing image-to-point cloud semantic mapping are widely used for traversable area recognition tasks [45, 51, 59], there is a lack of quantitative evaluation of their environmental perception results. Taking the ORD-BKI method as an example, a quantitative analysis of its perception performance is conducted by setting the parameter $k$ in the Bayesian Kernel Inference to 8. To assess the effect of the Bayesian Kernel Inference module, point clouds are voxelized using only semantic mapping, and evaluations are performed. The comparative results, presented in Table IV, demonstrate that Bayesian Kernel Inference significantly enhances voxel occupancy prediction. Nevertheless, since the inference relies solely on voxelized grids of sparse point clouds, all metrics of the ORD-BKI model remain substantially lower than those of the semantic scene completion-based approach.
TABLE IV: Impact of Bayesian Kernel Inference on Semantic Mapping.
@@
Notably, ORDformer achieves state-of-the-art results in the lethal and medium-cost categories, with IoU scores of 19.98 and 13.89, respectively. Specifically, in the lethal traversability category, ORDformer attains an IoU score of 19.98, which is 7.91 IoU higher than that of the second-best model, Monoscene. Furthermore, in the free and low-cost categories, ORDformer delivers competitive performance. The overall SSC mIoU and SC IoU scores are 15.70 and 49.78, respectively. These results represent significant improvements of over 20% in SC IoU compared to other methods, underscoring the excellent performance of ORDformer in occupancy prediction. Importantly, the enhanced geometric perception capabilities of ORDformer are critical for identifying complex negative and overhanging obstacles in off-road environments, where accurate 3D traversability estimation is crucial for safe navigation.
@@ IV-B Qualitative Analysis
Figure 8: Visualization of traversability prediction results from different models. From left to right: camera views, ground truth, predictions from the main algorithm (columns 3-6), and traversability estimations from ORDformer.”
Visualization of traversability prediction results from different models is presented in Fig. 8. By comparing these results with the ground truth, it is evident that ORDformer effectively identifies road traversability challenges in off-road environments. Additionally, the model accurately infers and completes regions containing vegetation and other obstacles.
Figure 9: Detection of rocky piles. Top images: camera views 10 seconds before (left), at the current time (center), and 10 seconds after (right). Bottom images: ground truth (left) and traversability estimations from ORDformer (right).
As illustrated in Fig. 9, a pile of rocks appears in front of the platform. Remarkably, the model effectively recognizes the rocky pile as impassable and accurately assesses its spatial distribution. Although the ground truth data designates the edges of the rocky obstacles as lethal regions, it incorrectly assigns labels to the flat areas within the rocky pile. Nevertheless, the model correctly infers the entire rocky pile as lethal, demonstrating its generalization capabilities.
Figure 10: Detection of multiple consecutive trees. Top images: camera views 10 seconds before (left), at the current time (center), and 10 seconds after (right). Bottom images: ground truth (left) and traversability estimations from ORDformer (right).
In off-road scenarios, trees are among the most common obstacles that ground unmanned platforms must avoid during trajectory planning. Therefore, accurate detection of trees is essential. Fig. 10 illustrates the detection results of ORDformer for multiple consecutively distributed trees. Different colored bounding boxes indicate the correspondence between various trees in images captured at different times, the ground truth, and traversability estimation results. By accurately identifying the spatial distribution of trees within the effective detection range, ORDformer ensures that ground unmanned platforms can effectively avoid these common obstacles in off-road environments, thereby facilitating safe trajectory planning.
@@
The inference results of the ORD-BKI model are visualized, as shown in Fig. 11. The results demonstrate that although the ORD-BKI model effectively identifies traversable ground areas in the two-dimensional plane, its performance in representing the 3D environment is suboptimal. This observation suggests that semantic mapping approaches are significantly influenced by the density of LiDAR point clouds. Such approaches are more suitable for platforms equipped with multiple supplementary blind-spot LiDARs. In contrast, unmanned platforms equipped with only a single LiDAR sensor experience limitations in representational performance for 3D traversable area recognition tasks due to the low point cloud density within the field of view of the camera.
Ground surfaces become partially exposed due to vehicle passage, resulting in the formation of ruts. These rutted areas typically exhibit sparse weed distribution and lack clear boundaries with the surrounding densely weed-covered regions. Their characteristics lie between dirt and grass, which poses significant challenges for accurate identification. Therefore, the recognition accuracy of ORDformer in these transitional areas requires further improvement.
@@ IV-C Real Vehicle Experiments
ORDformer is implemented on a real vehicle platform and validated through tests conducted at the Xiaotianshan Professional Outdoor Off-Road Site in Zhangjiakou City, Hebei Province, China. The test vehicle used is a Jeep Grand Cherokee, equipped with a Suteng Innovation 80-line LiDAR and a Senyun Intelligent monocular camera (see Fig. 12).
@@
The experiments encompass six typical scenarios: (a) rural roads without ground markings; (b) stepped obstacles composed of scattered rocks and earth mounds; (c) vehicle obstacles present in the direction of travel; (d) wide dirt roads with scattered weeds and shrubs; (e) dirt roads flanked by trees on both sides; and (f) muddy roads filled with deep vehicle ruts. These scenarios simulate diverse challenges in real off-road environments, aiming to evaluate the perception performance and robustness of the algorithm under varying complexities and dynamic conditions.
Figure 13: Real-vehicle test scenarios and model traversability estimation results: (a) Rural Path, (b) Stepped Barrier, (c) Vehicle Barrier, (d) Plain, (e) Forest, and (f) Mud Ruts.
@@
Despite these differences, the real-vehicle test results (see Fig. 13) demonstrate that the proposed algorithm accurately identifies the 3D traversability of terrains in off-road environments. This outcome indicates the robust adaptability and generalization capabilities of the algorithm, enabling effective operation across various sensor configurations and vehicle platforms. During the experiments, it is observed that under intense sunlight conditions (see Fig. 14), the input images exhibit purple halos and localized highlight overexposure effects, leading to incorrect recognitions by the model. Therefore, future work will involve enhancing the robustness of the model to drastic changes in environmental illumination to improve recognition accuracy under challenging lighting scenarios.

### v2

@@ III-C Architecture of ORD-BKI
The research presented in [54] introduces the S-BKI model, which infers grid attributes for discrete semantic grid maps to achieve continuous and dense semantic mapping. To ensure a fair comparison, the S-BKI model is constructed using LiDAR point clouds as input. Based on the vehicle’s geometric passability, a feasible domain recognition model, ORD-BKI, is developed, which employs Bayesian Kernel Inference for semantic scene completion, as illustrated in Fig. 4.
@@
The image semantic segmentation results are projected onto the point clouds to obtain semantically enriched point clouds. Due to the discrete and sparse nature of LiDAR point clouds, it is essential to model the distribution of grid semantic attributes during voxelization to infer the semantic attributes of neighboring grids [54]. The ORD-BKI model exploits Bayesian Kernel Inference to achieve semantic scene completion, approximating a Gaussian process, which reduces the time complexity from cubic to linear complexity [56]. The Nearest Neighbor algorithm is then used to assign semantics, generating dense semantic occupancy results. The model utilizes an anisotropic kernel function for inference and analyzes the model results under historical frame point cloud inputs. The geometric passability analysis method adopted by the model is consistent with the discussions demonstrated earlier in the article.
@@ IV-A Quantitative Analysis
Experiments are conducted on the RELLIS-OCC dataset for the tasks of 3D traversability cost estimation, of which the results are compared to those obtained by the algorithms presented in [57],[58],[59],[54]. Due to GPU memory constraints, the MonoScene model is trained on $4\times V100\ GPUs$, while the other models are trained on $4\times 3090\ GPUs$. A consistent batch size of 4 is maintained, and each model is trained for 20 epochs.
Evaluation Metrics: The RELLIS-OCC dataset comprises both semantic labels and traversability cost annotations. The traversability cost estimation task is framed as a semantic occupancy prediction task, encompassing four categories: lethal, medium-cost, low-cost, and free. Consequently, traditional metrics employed in 3D semantic occupancy prediction tasks are utilized for evaluation. As regards the scene completion (SC) task, the Intersection over Union (IoU) of occupied voxels, disregarding their semantic classes, is used as the evaluation metric. Correspondingly, with respect to the semantic scene completion (SSC) task, the IoU of each semantic class and the mean IoU (mIoU) across all semantic classes are adopted as evaluation metrics.
3D Traversability Cost Estimation:The results of the 3D Traversability Cost Estimation task are illustrated in Table II. Of all the methods, those on basis of semantic scene completion outperform the ORD-BKI method, which relies on Bayesian Kernel Inference. This emphasizes the efficiency of semantic scene completion approaches in enhancing off-road traversable area recognition. Notably, methods using only LiDAR point clouds as input, such as LMSCNet and SSCNet, exhibit inferior performance in the traversability cost estimation task compared to models that incorporate camera input. This suggests that image input is indispensable for accurate traversability cost estimation.
TABLE II: Traversability Estimate Results on the RELLIS-OCC Dataset.
@@
Lethal Medium-cost Low-cost Free mIoU IoU
@@
3DTTNet(Ours) 19.98 13.89 23.50 5.41 15.70 49.78
Our investigation reveals that while methods employing image-to-point cloud semantic mapping are widely used for traversable area recognition tasks [36],[42],[54], there is a lack of quantitative evaluation of their environmental perception results. Taking the ORD-BKI method as an example, a quantitative analysis of its perception performance is conducted by setting the parameter $k$ in the Bayesian Kernel Inference to 8. To assess the effect of the Bayesian Kernel Inference module, point clouds are voxelized using only semantic mapping, and evaluations are performed. The comparative results, presented in Table III, demonstrate that Bayesian Kernel Inference can greatly strengthen voxel occupancy prediction. Nevertheless, since the inference relies solely on voxelized grids of sparse point clouds, all metrics of the ORD-BKI model remain substantially lower than those of the semantic scene completion-based approach.
TABLE III: Impact of Bayesian Kernel Inference on Semantic Mapping.
@@
Notably, 3DTTNet achieves state-of-the-art results in the lethal and medium-cost categories, with IoU scores of 19.98 and 13.89, respectively. Specifically, in the lethal traversability category, 3DTTNet attains an IoU score of 19.98, which is 7.91 IoU higher than that of the second-best model, MonoScene. Furthermore, in the free and low-cost categories, 3DTTNet delivers competitive performance. The overall SSC mIoU and SC IoU scores are 15.70 and 49.78, respectively. These results represent significant improvements of 42% in SC IoU compared to other methods, indicating the excellent performance of 3DTTNet in occupancy prediction. Importantly, the enhanced geometric perception capabilities of 3DTTNet are critical for identifying complex negative and overhanging obstacles in off-road environments, where accurate 3D traversability estimation is crucial for safe navigation.
@@ IV-B Qualitative Analysis
Visualization of traversable terrain modeling results from different models is presented in Fig. 6. By comparing these results with the ground truth, it is evident that 3DTTNet effectively identifies road traversability challenges in off-road environments. Additionally, the model accurately infers and completes regions covering with vegetation and other obstacles.
Camera View
Semantic Annotations
Traversability Annotations
3DTTNet(Ours)
MonoScene
SSCNet
SSCNet-full
LMSCNet
LMSCNet-$S$
ORD-BKI
Figure 6: Visualization of traversability prediction results from various methods. The rows present different views and predictions, from top to bottom: (row 1) Camera views depicting real-world scenes, (row 2) Semantic annotations of the environment, (row 3) Traversability annotations representing the ground truth, (row 4) Traversability estimations produced by 3DTTNet, and (row 5) Predictions from alternative algorithms (rows 5-10). Color-category mappings are defined in Fig. 3(a) for semantic annotations and Fig. 3(b) for traversability annotations respectively.
As illustrated in Fig. 7, a pile of rocks appears in front of the platform. Remarkably, the model effectively recognizes the rocky pile as impassable and accurately assesses its spatial distribution. Although the ground truth data designates the edges of the rocky obstacles as lethal regions, it incorrectly assigns labels to the flat areas within the rocky pile. In contrast, 3DTTNet correctly infers the entire rocky pile as lethal, demonstrating its generalization capabilities.
Figure 7: Detection of rocky piles. Top images: camera views 10 seconds before (left), at the current time (center), and 10 seconds after (right). Bottom images: ground truth (left) and traversability estimations from 3DTTNet (right).
Figure 8: Detection of multiple consecutive trees. Top images: camera views 10 seconds before (left), at the current time (center), and 10 seconds after (right). Bottom images: ground truth (left) and traversability estimations from 3DTTNet (right).
@@
In off-road scenarios, trees are among the most common obstacles that ground unmanned platforms must avoid during trajectory planning. Therefore, accurate detection of trees is essential. Fig. 8 illustrates the detection results of 3DTTNet for multiple consecutively distributed trees. Different colored bounding boxes indicate the correspondence between various trees in images captured at different times, the ground truth, and traversability estimation results. By accurately identifying the spatial distribution of trees within the effective detection range, 3DTTNet enables ground unmanned platforms to effectively avoid these common obstacles in off-road environments, thereby facilitating safe trajectory planning.
The inference results of the ORD-BKI model are visualized, as shown in Fig. 9. The results demonstrate that although the ORD-BKI model effectively identifies traversable ground areas in the two-dimensional plane, its performance in representing the 3D environment is suboptimal. This observation suggests that semantic mapping approaches are significantly influenced by the density of LiDAR point clouds. Such approaches are more suitable for platforms equipped with multiple supplementary blind-spot LiDARs. In contrast, unmanned platforms equipped with only a single LiDAR sensor experience limitations in representational performance for 3D traversable area recognition tasks, due to the low point cloud density within the field of view of the camera.
Ground surfaces become partially exposed owing to vehicle passage, resulting in the formation of ruts. These rutted areas typically exhibit sparse weed distribution and lack clear boundaries with the surrounding densely weed-covered regions. Their characteristics lie between dirt and grass, which poses significant challenges for precise identification. Therefore, the recognition accuracy of 3DTTNet in these transitional areas requires further improvement.
@@ IV-C Real Vehicle Experiments
3DTTNet is implemented on a real vehicle platform and validated through tests conducted at the Xiaotianshan Professional Outdoor Off-Road Site in Zhangjiakou City, Hebei Province, China. The test vehicle is a Jeep Grand Cherokee, equipped with a Suteng Innovation 80-line LiDAR and a Senyun Intelligent monocular camera (see Fig. 10).
@@ IV-C1 Test 1
The experiments are involved with six typical scenarios: (a) Rural path: rural roads without ground markings; (b) Stepped barrier: stepped obstacles composed of scattered rocks and earth mounds; (c) Vehicle barrier: vehicle obstacles presented in the direction of travel; (d) Plain: wide dirt roads with scattered weeds and shrubs; (e) Forest: dirt roads flanked by trees on both sides; and (f) Mud ruts: muddy roads filled with deep vehicle ruts. These scenarios simulate diverse challenges in real off-road environments, aiming to evaluate the perception performance and robustness of the algorithm under varying complexities and dynamic conditions.
Results
(a) Rural
Path
(b) Stepped
Barrier
(c) Vehicle
Barrier
(d) Plain
(e) Forest
(f) Mud Ruts
Camera View
3DTTNet(Ours)
MonoScene
SSCNet
SSCNet-full
LMSCNet
LMSCNet-$S$
ORD-BKI
Figure 11: Real-vehicle test scenarios conducted at the Xiaotianshan Professional Outdoor Off-Road Site. The six test scenarios, presented from left to right, are: (a) Rural Path, (b) Stepped Barrier, (c) Vehicle Barrier, (d) Plain, (e) Forest, and (f) Mud Ruts, along with traversability estimation results from different algorithms. The first row displays the camera views, followed by the estimations from 3DTTNet (Ours), MonoScene, SSCNet, SSCNet-full, LMSCNet, LMSCNet-$S$, and ORD-BKI. Different regions are labeled according to their traversability, with color-category mapping detailed in Fig. 3(b).
TABLE IV: Traversability estimate results from the real vehicle test. The table summarizes the inference times and performance across six test scenarios (labeled (a) to (f)) for various methods. Methods that successfully estimated traversability in each scenario are marked with a check ($\checkmark$), while methods that did not perform the estimation are marked with a dash (–).
Method Inference Time Scenario (a) Scenario (b) Scenario (c) Scenario (d) Scenario (e) Scenario (f)
MonoScene[57] 4410ms - - - - - -
SSCNet[58] 31ms $\checkmark$ - - $\checkmark$ - -
SSCNet-full[58] 24ms $\checkmark$ - - - - -
LMSCNet[59] 63ms - - - - - -
LMSCNet-SS[59] 53ms $\checkmark$ - - $\checkmark$ $\checkmark$ -
ORD-BKI[54] 5243ms - - - - - -
3DTTNet(Ours) 340ms $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Despite these differences, the real-vehicle test results (see Fig. 11 and Table IV) demonstrate that the proposed algorithm accurately identifies the 3D traversability of terrains in off-road environments. This outcome indicates the robust adaptability and generalization capabilities of the algorithm, enabling effective operation across various sensor configurations and vehicle platforms.
IV-C2 Test 2
(a) Cliff Terrain
(b) Constrained Forest Road
Camera View
Trajectory
Camera View
Trajectory
(c) Continuous Irregular Pits
(d) Water-filled Ruts
Camera View
Trajectory
Camera View
Trajectory
Figure 12: Multi-scenario trajectory planning visualization in off-road environments. Experimental results for four representative scenarios: (a) Cliff terrain, (b) Constrained forest road, (c) Continuous irregular pits, and (d) Water-filled ruts. Each case demonstrates three consecutive input frames (left panels, arranged chronologically from top to bottom) and corresponding planning outputs (right panels). The Hybrid A* algorithm generates initial trajectory nodes (light pink grids), while the Reeds-Shepp algorithm produces optimized smooth trajectories (dark blue curves).
We further validated the practicality of the proposed framework through trajectory planning experiments. Leveraging the traversability cost map generated by 3DTTNet, we integrated vehicle kinematic constraints into the Hybrid A* algorithm to search for collision-free initial trajectories in complex off-road environments. To ensure dynamic feasibility, the Reeds-Shepp algorithm was subsequently applied to refine trajectory smoothness and adherence to motion constraints. The experiments encompass representative scenarios, including: (a) Cliff terrain: A highly unstructured environment with a cliff on the road’s left side, loose rock piles along the edge, and a steep slope beneath, challenging both perception and trajectory safety; (b) Constrained forest road: A narrow, winding path flanked by dense vegetation, requiring precise navigation in confined spaces with limited visibility; (c) Continuous irregular pits: A forest road with densely distributed irregular pits on the right side, severely degrading vehicle passability and demanding adaptive obstacle avoidance; (d) Water-filled ruts: Deep ruts formed by preceding vehicles, compounded by rainwater accumulation, introducing slippery surfaces and unpredictable terrain deformation.
These scenarios simulate extreme yet realistic challenges in off-road environments, including non-uniform obstacles, dynamic terrain conditions, and constrained mobility. As demonstrated in Fig. 12, our framework generates real-time trajectories that balance safety and kinematic feasibility across these scenarios. The results highlight the system’s robustness in heterogeneous environments and its potential to address critical gaps in autonomous off-road navigation, such as handling irregular geometries and partial occlusions.
From the experiments, it is observed that under intense sunlight conditions (see Fig. 13), the input images exhibit purple halos and localized highlight overexposure effects, leading to incorrect recognitions of the model. Accordingly, future work will involve enhancing the robustness of the model to drastic changes in environmental illumination to improve recognition accuracy under challenging lighting scenarios.